# FlashTex: Fast Relightable Mesh Texturing with LightControlNet

### v1

@@ Abstract
Manually creating textures for 3D meshes is time-consuming, even for expert visual content creators. We propose a fast approach for automatically texturing an input 3D mesh based on a user-provided text prompt. Importantly, our approach disentangles lighting from surface material/reflectance in the resulting texture so that the mesh can be properly relit and rendered in any lighting environment. We introduce LightControlNet, a new text-to-image model based on the ControlNet architecture, which allows the specification of the desired lighting as a conditioning image to the model. Our text-to-texture pipeline then constructs the texture in two stages. The first stage produces a sparse set of visually consistent reference views of the mesh using LightControlNet. The second stage applies a texture optimization based on Score Distillation Sampling (SDS) that works with LightControlNet to increase the texture quality while disentangling surface material from lighting. Our pipeline is significantly faster than previous text-to-texture methods, while producing high-quality and relightable textures. Project page: https://flashtex.github.io/.
{strip}
Figure 1: We propose an efficient approach for texturing an input 3D mesh given a user-provided text prompt. Notably, our generated texture can be relit properly in different lighting environments. We show two viewpoints rendered with two different lighting conditions.
@@ 1 Introduction
Creating high-quality textures for 3D meshes is crucial across industries such as gaming, film, animation, AR/VR, and industrial design. Traditional mesh texturing tools are labor-intensive, and require extensive training in visual design. As the demand for immersive 3D content continues to surge, there is a pressing need to streamline and automate the mesh texturing process.
@@
Figure 2: Given a 3D mesh of a helmet (a) and a lighting environment $L$, the reference rendering (b) depicts the “correct” highlights on the mesh due to $L$, by treating its surface reflectance as half-metal and half-smooth with a gray diffuse color. (c) The texture generated by Fantasia3D [8] is not properly relit for the lighting $L$ because Fantasia3D bakes most of the lighting into the diffuse texture for the mesh and does not capture the bright highlights in the specular texture. (d) In contrast, our text-to-texture pipeline disentangles lighting from material, better capturing the diffuse and specular components of the metal helmet in this environment.
In this work, we propose an efficient approach for texturing an input 3D mesh based on a user-provided text prompt that disentangles the lighting from surface material/reflectance to enable relighting (Figure FlashTex: Fast Relightable Mesh Texturing with LightControlNet). Our method introduces LightControlNet, an illumination-aware text-to-image diffusion model based on the ControlNet [57] architecture, which allows specification of the desired lighting as a conditioning image for the diffusion model. Our text-to-texture pipeline uses LightControlNet to generate relightable textures in two stages. In stage 1, we use multi-view visual prompting in combination with the LightControlNet to produce visually consistent reference views of the 3D mesh for a small set of viewpoints. In stage 2, we perform a new texture optimization procedure that uses the reference views from stage 1 as guidance, and extends Score Distillation Sampling (SDS) [35] to work with LightControlNet. This allows us to increase the texture quality while disentangling the lighting from surface material/reflectance. We show that the guidance from the reference views allows our optimization to generate textures with over 10x speed-up than previous SDS-based relightable texture generation methods such as Fantasia3D [8]. Furthermore, our experiments show that the quality of our textures is generally better than those of previous baseline methods in terms of FID, KID, and user evaluations.
@@ 2 Related Work
Text-to-Image Generation. Recent years have seen significant advancements in text-to-image generation empowered by diffusion models [37, 40, 39]. Stable Diffusion [39], for example, trains a latent diffusion model (LDM) on the latent space rather than pixel space, delivering highly impressive results with relatively affordable computational costs. Further extending the scope of text-based diffusion models, works such as GLIGEN [19], PITI [50], T2IAdapter [27], and ControlNet [57] incorporate spatial conditioning inputs (e.g., depth maps, normal maps, edge maps, etc.) to enable localized control over the composition of the result. Beyond their power in image generation, these 2D diffusion models, trained on large-scale text-image paired datasets, also contribute valuable priors to various other tasks such as image editing [13, 24], recognition [54], and 3D generation [35].
Text-to-3D Generation. The success of text-to-image generation has sparked considerable interest in its 3D counterpart. Some approaches [43, 59, 17, 30] train a text-conditioned 3D generative model akin to 2D models, while others employ 2D priors from pre-trained diffusion models for optimization [35, 49, 21, 8, 25, 52, 45, 18] and multi-view synthesis [23, 42]. For instance, DreamFusion [35] and Score Jacobian Chaining [49] were the first to propose Score Distillation Sampling to optimize a 3D representation using 2D diffusion model gradients. Zero-1-to-3 [23] synthesizes novel views using a pose-conditioned 2D diffusion model. Yet, these methods often produce blurry, low-frequency textures that bake lighting into surface reflectance. Fantasia3D [8] can generate more realistic textures by incorporating physics-based materials. However, the resulting materials remain entangled with lighting, making it difficult to relight the textured object in a new lighting environment. In contrast, our method can effectively disentangle the lighting and surface reflectance texture. Concurrent to our work, MATLABER [55] aims to recover material information in text-to-3D generation using a material autoencoder. Our method, however, differs in approach and improves efficiency.
@@
Closest to our work are the recent works that use pre-trained 2D diffusion models and treat texture generation as a byproduct of text-to-3D generation. Examples include Latent-Paint [25], which uses Score Distillation Sampling in latent space, Text2tex [7], which leverages depth-based 2D ControlNet, and TEXTure [38], which exploits both previous methods. Nonetheless, similar to recent text-to-3D models, such methods produce textures with entangled lighting effects and suffer from slow generation. On the other hand, TANGO [9], generates material textures using a Spherical-Gaussian-based differentiable renderer, but struggles with complex texture generation.
@@ 3 Preliminaries
Score Distillation Sampling (SDS). DreamFusion [35] optimizes a 3D scene representation conditioned on text prompts using a pre-trained 2D text-to-image diffusion model. The scene is represented as a NeRF  [26, 1] parametrization $\theta$. A differentiable renderer $\mathcal{R}$ applied to $\theta$ with a randomly sampled camera view $C$ then generates a 2D image $x=\mathcal{R}(\theta,C)$. A small amount of noise $\epsilon\sim\mathbb{N}(0,1)$ is then added to $x$ to obtain a noisy image $x_{t}$. DreamFusion leverages a diffusion model $\phi$ (Imagen [40]) to provide a score function $\hat{\epsilon}_{\phi}(x_{t};y,t)$, which predicts the sampled noise $\epsilon$ given the noisy image $x_{t}$, text prompt $y$, and noise level $t$. This score function guides the direction of the gradient for updating the scene parameters $\theta$, and the gradient is calculated by Score Distillation Sampling (SDS):
@@ 4 Method
Our text-to-texture pipeline operates in two main stages to generate a relightable texture for an input 3D mesh with a corresponding text prompt (Figure 3). In stage 1, we use a multi-view visual prompting approach to obtain visually consistent views of the object from a small set of viewpoints, using a 2D ControlNet. Simply backprojecting these sparse views onto the 3D mesh could produce patches of high-quality texture, but would also generate visible seams and other visual artifacts where the views do not fully match. The resulting texture would also have lighting baked-in, making it difficult to relight the textured mesh in a new lighting environment. Therefore, in stage 2, we apply a texture optimization that uses a ContolNet in combination with Score Distillation Sampling (SDS) [35] to mitigate such artifacts and separate lighting from the surface material properties/reflectance. In both stages, we introduce a new illumination-aware ControlNet that allows us to specify the desired lighting as a conditioning image for an underlying text-to-image diffusion model. We call this model LightControlNet and describe how it works in Section Section 4.1. We then detail each stage in Sections Section 4.2 and Section 4.3, respectively.
Figure 4: (a) LightControlNet requires a conditioning image that specifies desired lighting $L$ for a view $C$ of a 3D mesh. To form the conditioning image, we first render the mesh with the desired $L$ and $C$ using three different materials: (1) non-metal, not smooth, (2) half-metal, half-smooth, and (3) pure metal, smooth, and then combine the renderings into a single three-channel image. (b) LightControlNet is a diffusion model that takes such light conditioning images as input as well as a text prompt and generates images accordingly.
@@ 4.1 LightControlNet
LightControlNet adapts the ControlNet architecture to enable control over the lighting in the generated image. More specifically, we create a conditioning image for a 3D mesh by rendering it using three pre-defined materials and under known lighting conditions (Figure 4). These renderings encapsulate information about the desired shape and lighting for the object, and we stack them into a three-channel conditioning image. We have found that setting the pre-defined materials to (1) non-metal, non-smooth; (2) non-metal, half smooth; and (3) pure metal, extremely smooth, respectively, works well in practice.
@@
To train our LightControlNet, we use 40K objects from the Objaverse dataset [10]. Each object is rendered from 12 views using a randomly sampled camera $C$ and lighting $L$ sampled from 6 environment maps sourced from the Internet. For each resulting $(L,C)$ pair, we render the conditioning image using the pre-defined materials, as well as the full-color rendering of the object using its original materials and textures. We use the resulting 480K pairs of (conditioning images, full-color rendering) to train LightControlNet using the approach of Zhang et al. [57].
@@
Once LightControlNet is trained, we can specify the desired view and lighting for any input 3D mesh. We first render the conditioning image with the desired view and lighting and then pass it along with a text prompt into LightControlNet, to obtain high-quality images. These images are spatially aligned to the desired view, lit with the desired lighting, and contain detailed surface textures (Figure FlashTex: Fast Relightable Mesh Texturing with LightControlNet).
@@
Distilling the encoder. We improve the efficiency of LightControlNet by distilling the image encoder in Stable Diffusion [39], the base diffusion model in the ControlNet architecture. The original Stable Diffusion image encoder consumes almost 50% of the forward and backward time of SDS calculation using the latent diffusion model, primarily in downsampling the input image. Metzer et al. [25] have found the image decoder from latent space to image space can be closely approximated by per-pixel matrix multiplication. Inspired by this, we distill the encoder by removing its attention modules and training it on the COCO dataset [22] to match the original output. This distilled encoder runs 5x faster than the original one, resulting in an approximately 2x acceleration of our text-to-texture pipeline without compromising output quality. An ablation study of our distilled encoder is detailed in Table 2, with additional implementation specifics in Section A of the appendix.
@@
Figure 5: Multi-View Visual Prompting. (a) When we independently input four canonical conditioning images to LightControlNet, it generates four very different appearances and styles even with a fixed random seed. (b) When we concatenate the four images into a 2$\times$2 grid and pass them as a single image into LightControlNet, it produces a far more consistent appearance and style. Text prompt: “Hiking Boot”.
@@ 4.2 Stage 1: Multi-view Visual Prompting
In stage 1, we leverage LightControlNet to synthesize high-quality 2D images for a sparse set of views of the 3D mesh. Specifically, we create conditioning images for four canonical views $C^{*}$ around the equator of the 3D mesh using a fixed lighting environment map $L^{*}$. One approach to generating the complete texture for the mesh would be to apply the LightControlNet independently with each such conditioning image, but using the same text prompt, and then backprojecting the four output images to the surface of the 3D mesh. In practice, however, applying the LightControlNet to each view independently produces inconsistent images of varying appearance and style, even when the text prompt and random seed remain fixed (Figure 5).
To mitigate this multi-view inconsistency issue, we take a multi-view visual prompting approach. We concatenate the conditioning images for the four canonical views into a single 2 $\times$ 2 grid and treat it as a single conditioning image. We observe that applying LightControlNet to all four views simultaneously, using this combined multi-view conditioning image, results in a far more consistent appearance and style across the views, compared to independent prompting (Figure 5). We suspect this property arises from the presence of similar training data samples – grid-organized sets depicting the same object – in Stable Diffusion’s training set, which is also observed in concurrent works [53, 58]. Formally, we generate the conditioning image $I_{\text{cond}}(L^{*},C^{*})$ under a fixed canonical lighting condition $L^{*}$ using the four canonical viewpoints $C^{*}$. We then apply our LightControlNet with text prompt $y$, to generate the corresponding reference output image $I_{\text{ref}}$ as
@@ 4.3 Stage 2: Texture Optimization
To address both of these issues, we employ a texture optimization using SDS loss. Specifically, we use a multi-resolution hash-grid [28] as our 3D scene representation, instead of NeRF as in the original DreamFusion formulation [35]. Given a 3D point $p\in\mathbb{R}^{3}$ on the mesh, our hash-grid produces a 32-dimensional multi-resolution feature. This feature is then fed to a 2-layer MLP $\Gamma$ to obtain the texture material parameters for this point. Similar to Fantasia3D [8], these material parameters consist of metallicness $k_{m}\in\mathbb{R}$, roughness $k_{r}\in\mathbb{R}$, a bump vector $k_{n}\in\mathbb{R}^{3}$ and the base color $k_{c}\in\mathbb{R}^{3}$. Formally,
@@
where $\beta$ is the multi-resolution hash encoding function. Notably, this 3D hash-grid representation can be easily converted to 2D uv texture maps, which are more friendly to downstream applications. Given the mesh $M$, the texture $\Gamma(\beta(\cdot))$, a camera view $C$ and lighting $L$ we can use nvdiffrast [16], a differentiable renderer $\mathcal{R}$ to produce a 2D rendering of it $x$, as
@@
Recall that the optimization approach of DreamFusion [35] randomly samples camera views $C$, generates an image for $C$ using diffusion model $\phi$, and supervises the optimization using the SDS loss. We extend this optimization in two ways. First, we use four fixed reference images $I_{\text{ref}}$ with its canonical camera views $C^{*}$ and lighting $L^{*}$ to guide the texture optimization, through a reconstruction loss that we apply whenever we sample a canonical view.
$\mathcal{L}_{\text{recon}}$ $=\lambda_{\text{MSE}}\mathcal{L}_{\text{MSE}}[I_{\text{ref}},\mathcal{R}(\Gamma(\beta(\cdot)),L^{*},C^{*})]$
$+\lambda_{\text{VGG}}\mathcal{L}_{\text{VGG}}[I_{\text{ref}},\mathcal{R}(\Gamma(\beta(\cdot)),L^{*},C^{*})].$
When we sample a non-canonical view $C$, we sample a random lighting $L$ and use the SDS loss to supervise the optimization, but with our LightControlNet as the diffusion model $\phi_{\text{LCN}}$, so
$\nabla_{\Gamma,\beta}\mathcal{L}_{\text{SDS}}(\phi_{\text{LCN}},x)$
$=$ $\mathbb{E}_{t,\epsilon}\left[w(t)(\hat{\epsilon}_{\phi_{\text{LCN}}}(x_{t};y,t,I_{\text{cond}}(L,C))-\epsilon)\frac{\partial x}{\partial\Gamma(\beta(\cdot))}\right],$
@@
where $S$ denotes the object surface and $\epsilon$ is a small random 3D perturbation.
Scheduling the optimization. We warm up the optimization by rendering the 4 canonical views and applying $\mathcal{L}_{\text{recon}}$ for 50 iterations. We then add in iterations using the $\mathcal{L}_{\text{SDS}}$ loss and optimize over randomly chosen camera views and randomly selected lighting from a pre-defined set of environmental lighting maps. Specifically we alternate iterations between using $\mathcal{L}_{\text{SDS}}$ and $\mathcal{L}_{\text{recon}}$. In addition, for a quarter of the SDS iterations, we use the canonical views rather than randomly selecting the views. This ensures that the resulting texture does not overfit to the reference images corresponding to the canonical views. The warm-up iterations capture the large-scale structure of our texture and allow us to use relatively small noise levels ($t\leq 0.1$) in the SDS optimization. We sample the noise following a linearly decreasing schedule [15] with $t_{\text{max}}=0.1$ and $t_{\text{min}}=0.02$. We also adjust the conditioning strength $s$ of our LightControlNet in the SDS loss linearly from $1$ to 00 over these iterations so that LightControlNet is only lightly applied by the end of the optimization. We have experimentally found that we obtain high-quality textures after 400 total iterations of this optimization and this is far fewer iterations than other SDS-based texture generation techniques such as Fantasia3D [8] which requires 5000 iterations. More implementation details are in Section A of the appendix.
Faster pipeline with baked-in lighting. Our text-to-texture pipeline primarily uses the lighting control capabilities of LightControlNet in the texture optimization process to separate lighting from surface reflectance. We have experimented with replacing the LightControlNet in stage 1 of our pipeline with a depth ControlNet that uses a depth rendering of the mesh as the conditioning image and replacing it with Stable Diffusion [39] based SDS in stage 2. While this approach leaves lighting baked into the resulting texture, it does increase the speed of our pipeline, because it eliminates the forward pass of LightControlNet in the SDS optimization. As shown in Table 1, we gain 2$\times$ speedup with this version.

### v2

@@ Abstract
Manually creating textures for 3D meshes is time-consuming, even for expert visual content creators. We propose a fast approach for automatically texturing an input 3D mesh based on a user-provided text prompt. Importantly, our approach disentangles lighting from surface material/reflectance in the resulting texture so that the mesh can be properly relit and rendered in any lighting environment. We introduce LightControlNet, a new text-to-image model based on the ControlNet architecture, which allows the specification of the desired lighting as a conditioning image to the model. Our text-to-texture pipeline then constructs the texture in two stages. The first stage produces a sparse set of visually consistent reference views of the mesh using LightControlNet. The second stage applies a texture optimization based on Score Distillation Sampling (SDS) that works with LightControlNet to increase the texture quality while disentangling surface material from lighting. Our algorithm is significantly faster than previous text-to-texture methods, while producing high-quality and relightable textures.
@@ 1 Introduction
Creating high-quality textures for 3D meshes is crucial across industries such as gaming, film, animation, AR/VR, and industrial design. Traditional mesh texturing tools are labor-intensive, and require extensive training in visual design. As the demand for immersive 3D content continues to surge, there is a pressing need to streamline and automate the mesh texturing process (Figure 1).
@@
Figure 1: We propose an efficient approach for texturing an input 3D mesh given a user-provided text prompt. Our generated texture can be relit properly in different lighting environments. The light probe shows the varied lighting environment. We suggest the readers check our video results of rotating lighting in our supplementary material.
In this work, we propose an efficient approach for texturing an input 3D mesh based on a user-provided text prompt that disentangles the lighting from surface material/reflectance to enable relighting (Figure 1). Our method introduces LightControlNet, an illumination-aware text-to-image diffusion model based on the ControlNet [63] architecture, which allows specification of the desired lighting as a conditioning image for the diffusion model. Our text-to-texture pipeline uses LightControlNet to generate relightable textures in two stages. In stage 1, we use multi-view visual prompting in combination with the LightControlNet to produce visually consistent reference views of the 3D mesh for a small set of viewpoints. In stage 2, we perform a new texture optimization procedure that uses the reference views from stage 1 as guidance, and extends Score Distillation Sampling (SDS) [39] to work with LightControlNet. This allows us to increase the texture quality while disentangling the lighting from surface material/reflectance. We show that the guidance from the reference views allows our optimization to generate textures with over 10x speed-up than previous SDS-based relightable texture generation methods such as Fantasia3D [9]. Furthermore, our experiments show that the quality of our textures is generally better than those of existing baselines in terms of FID, KID, and user study.
Figure 2: Given a 3D mesh of a helmet (a) and a lighting environment $L$, the reference rendering (b) depicts the “correct” highlights on the mesh due to $L$, by treating its surface reflectance as half-metal and half-smooth with a gray diffuse color. (c) The texture generated by the leading method Fantasia3D [9] is not properly relit as Fantasia3D bakes most of the lighting into the diffuse texture for the mesh and does not capture the bright highlights in the specular texture. (d) In contrast, our pipeline disentangles lighting from material, better capturing the diffuse and specular components of the metal helmet in this environment. Text prompt: “A medieval steel helmet.”
@@ 2 Related Work
Text-to-Image generation. Recent years have seen significant advancements in text-to-image generation empowered by diffusion models [42, 45, 44]. Stable Diffusion [44], for example, trains a latent diffusion model (LDM) on the latent space rather than pixel space, delivering highly impressive results with affordable computational costs. Further extending the scope of text-based diffusion models, works such as GLIGEN [23], PITI [56], T2IAdapter [31], and ControlNet [63] incorporate spatial conditioning inputs (e.g., depth maps, normal maps, edge maps, etc.) to enable localized control over the composition of the result. Beyond their power in image generation, these 2D diffusion models, trained on large-scale text-image paired datasets, also contribute valuable priors to various other tasks such as image editing [15, 28], 3D generation [39, 41], and 3D editing [13, 19, 54, 66].
Text-to-3D synthesis. The success of text-to-image synthesis has sparked considerable interest in its 3D counterpart. Some approaches [48, 65, 21, 34] train a text-conditioned 3D generative model akin to 2D models, while others employ 2D priors from pre-trained diffusion models for optimization [39, 55, 25, 9, 29, 57, 50, 22] and multi-view synthesis [27, 47]. For instance, DreamFusion [39] and Score Jacobian Chaining [55] were the first to propose Score Distillation Sampling to optimize a 3D representation using 2D diffusion model gradients. Zero-1-to-3 [27] synthesizes novel views using a pose-conditioned 2D diffusion model. Yet, these methods often produce blurry, low-frequency textures that bake lighting into surface reflectance. Fantasia3D [9] can generate more realistic textures by incorporating physics-based materials. However, the resulting materials remain entangled with lighting, making it difficult to relight the textured object in a new lighting environment. In contrast, our method effectively disentangles the lighting and surface reflectance texture. Concurrent to our work, MATLABER [60] aims to recover material information in text-to-3D generation using a material autoencoder. Our method, however, differs in approach and improves efficiency.
@@
Closest to our work are the recent works that use pre-trained 2D diffusion models and treat texture generation as a byproduct of text-to-3D generation. Examples include Latent-Paint [29], which uses Score Distillation Sampling in latent space, Text2tex [8], which leverages depth-based 2D ControlNet, and TEXTure [43], which exploits both previous methods. Nonetheless, similar to recent text-to-3D models, such methods produce textures with entangled lighting effects and suffer from slow generation. On the other hand, TANGO [10], generates material textures using a Spherical-Gaussian-based differentiable renderer, but struggles with complex texture generation. A concurrent work, Paint3D [62], aims to generate lighting-less textures, yet it cannot produce material-based textures like ours.
@@ 3 Preliminaries
Score Distillation Sampling (SDS). DreamFusion [39] optimizes a 3D scene representation conditioned on text prompts using a pre-trained 2D text-to-image diffusion model. The scene is represented as a NeRF  [30, 1] parametrization $\theta$. A differentiable renderer $\mathcal{R}$ applied to $\theta$ with a randomly sampled camera view $C$ then generates a 2D image $x=\mathcal{R}(\theta,C)$. A small amount of noise $\epsilon\sim\mathbb{N}(0,1)$ is then added to $x$ to obtain a noisy image $x_{t}$. DreamFusion leverages a diffusion model $\phi$ (Imagen [45]) to provide a score function $\hat{\epsilon}_{\phi}(x_{t};y,t)$, which predicts the sampled noise $\epsilon$ given the noisy image $x_{t}$, text prompt $y$, and noise level $t$. This score function can update the scene parameters $\theta$, using the gradient calculated by Score Distillation Sampling (SDS):
@@ 4 Method
Our text-to-texture pipeline operates in two main stages to generate a relightable texture for an input 3D mesh with a corresponding text prompt (Figure 3). In Stage 1, we use a multi-view visual prompting approach to obtain visually consistent views of the object from a small set of viewpoints, using a 2D ControlNet. Simply backprojecting these sparse views onto the 3D mesh could produce patches of high-quality texture, but would also generate visible seams and other visual artifacts where the views do not fully match. The resulting texture would also have lighting baked-in, making it difficult to relight the textured mesh in a new lighting environment. Therefore, in Stage 2, we apply a texture optimization that uses a ControlNet in combination with Score Distillation Sampling (SDS) [39] to mitigate such artifacts and separate lighting from the surface material properties/reflectance. In both stages, we introduce a new illumination-aware ControlNet that allows us to specify the desired lighting as a conditioning image for an underlying text-to-image diffusion model. We call this model LightControlNet and describe how it works in Section 4.1. We then detail each stage in Section 4.2 and Section 4.3, respectively.
Figure 4: (a) LightControlNet requires a conditioning image that specifies desired lighting $L$ for a view $C$ of a 3D mesh. To form the conditioning image, we first render the mesh with the desired $L$ and $C$ using three different materials: (1) non-metal, not smooth, (2) half-metal, half-smooth, and (3) pure metal, smooth, and then combine the renderings into a single three-channel image. (b) LightControlNet is a diffusion model that is conditional on such light conditioning images as well as text prompts.
@@ 4.1 LightControlNet
LightControlNet adapts the ControlNet architecture to enable control over the lighting in the generated image. More specifically, we create a conditioning image for a 3D mesh by rendering it using three pre-defined materials and under known lighting conditions (Figure 4). These renderings encapsulate information about the desired shape and lighting for the object, and we stack them into a three-channel conditioning image. We have found that setting the pre-defined materials to (1) non-metal, not smooth; (2) half-metal, half-smooth; and (3) pure metal, extremely smooth, respectively, works well in practice. The specific material parameters are in the appendix.
@@
To train our LightControlNet, we use 40K objects from the Objaverse dataset [11]. Each object is rendered from 12 views using a randomly sampled camera $C$ and lighting $L$ sampled from 6 environment maps sourced from the Internet. $L$ is also subject to random rotation and intensity scaling. For each resulting $(L,C)$ pair, we render the conditioning image using the pre-defined materials, as well as the full-color rendering of the object using its original materials and textures. We use the resulting 480K pairs of (conditioning images, full-color rendering) to train LightControlNet using the approach of Zhang et al. [63].
@@
Once LightControlNet is trained, we can specify the desired view and lighting for any 3D mesh. We first render the conditioning image with the desired view and lighting and then pass it along with a text prompt into LightControlNet, to obtain high-quality images. These images are spatially aligned to the desired view, lit with the desired lighting, and contain detailed textures (Figure 4).
@@
Distilling the encoder. We improve the efficiency of LightControlNet by distilling the image encoder in Stable Diffusion [44], the base diffusion model in the ControlNet architecture. The original Stable Diffusion image encoder consumes almost 50% of the forward and backward time of SDS calculation using the latent diffusion model, primarily in downsampling the input image. Metzer et al. [29] have found the image decoder from latent space to image space can be closely approximated by per-pixel matrix multiplication. Inspired by this, we distill the encoder by removing its attention modules and training it on the COCO dataset [26] to match the original output. This distilled encoder runs 5x faster than the original one, resulting in an approximately 2x acceleration of our text-to-texture pipeline without compromising output quality. An ablation study of our distilled encoder is detailed in Table 3, with additional implementation specifics in the appendix.
@@
Figure 5: Multi-view visual prompting. (a) When we independently input four canonical conditioning images to LightControlNet, it generates four very different appearances and styles even with a fixed random seed. (b) When we concatenate the four images into a 2$\times$2 grid and pass them as a single image into LightControlNet, it produces a far more consistent appearance and style. Text prompt: “A hiking boot”.
@@ 4.2 Stage 1: Multi-view Visual Prompting
In Stage 1, we leverage LightControlNet to synthesize high-quality 2D images for a sparse set of views of the 3D mesh. Specifically, we create conditioning images for four canonical views $C^{*}$ around the equator of the 3D mesh using a fixed lighting environment map $L^{*}$ sampled from a set of environment maps. One approach to generating the complete texture for the mesh would be to apply the LightControlNet independently with each such conditioning image, but using the same text prompt, and then backprojecting the four output images to the surface of the 3D mesh. In practice, however, applying the LightControlNet to each view independently produces inconsistent images of varying appearance and style, even when the text prompt and random seed remain fixed (Figure 5).
To mitigate this multi-view inconsistency issue, we take a multi-view visual prompting approach. We concatenate the conditioning images for the four canonical views into a single 2 $\times$ 2 grid and treat it as a single conditioning image. We observe that applying LightControlNet to all four views simultaneously, using this combined multi-view conditioning image, results in a far more consistent appearance and style across the views, compared to independent prompting (Figure 5). We suspect this property arises from the presence of similar training data samples – grid-organized sets depicting the same object – in Stable Diffusion’s training set, which is also observed in concurrent works [59, 64]. Formally, we generate the conditioning image $I_{\text{cond}}(L^{*},C^{*})$ under a fixed canonical lighting condition $L^{*}$ using four canonical viewpoints $C^{*}$. We then apply our LightControlNet with text prompt $y$ to generate the corresponding reference image $I_{\text{ref}}$:
@@ 4.3 Stage 2: Texture Optimization
To address both of these issues, we employ texture optimization using SDS loss. Specifically, we use a multi-resolution hash-grid [32] as our 3D scene representation, instead of NeRF as in the original DreamFusion formulation [39]. Given a 3D point $p\in\mathbb{R}^{3}$ on the mesh, our hash-grid produces a 32-dimensional multi-resolution feature. This feature is then fed to a 2-layer MLP $\Gamma$ to obtain the texture material parameters for this point. Similar to Fantasia3D [9], these material parameters consist of metallicness $k_{m}\in\mathbb{R}$, roughness $k_{r}\in\mathbb{R}$, a bump vector $k_{n}\in\mathbb{R}^{3}$ and the base color $k_{c}\in\mathbb{R}^{3}$. Formally,
@@
where $\beta$ is the multi-resolution hash encoding function. Notably, this 3D hash-grid representation can be easily converted to 2D uv texture maps, which are more friendly to downstream applications. Given the mesh $M$, the texture $\Gamma(\beta(\cdot))$, a camera view $C$ and lighting $L$ we can use nvdiffrast [20], a differentiable renderer $\mathcal{R}$ to produce a 2D rendering of it, $x$, as
@@
Recall that the optimization approach of DreamFusion [39] randomly samples camera views $C$, generates an image for $C$ using diffusion model $\phi$, and supervises the optimization using the SDS loss. We extend this optimization in two ways. First, we use four fixed reference images $I_{\text{ref}}$ with their canonical views $C^{*}$ and lighting $L^{*}$ to guide the texture optimization through a reconstruction loss:
$\mathcal{L}_{\text{recon}}=||I_{\text{ref}}-\mathcal{R}(\Gamma(\beta(\cdot)),L^{*},C^{*})||_{2}+\mathcal{L}_{\text{perceptual}}(I_{\text{ref}},\mathcal{R}(\Gamma(\beta(\cdot)),L^{*},C^{*})),$
where both L2 loss and perceptual loss [18] are used. For a non-canonical view $C$, we sample a random lighting $L$ and use the SDS loss to supervise the optimization, but with our LightControlNet as the diffusion model $\phi_{\text{LCN}}$, so
$\nabla_{\Gamma,\beta}\mathcal{L}_{\text{SDS}}(\phi_{\text{LCN}},x)=\mathbb{E}_{t,\epsilon}\left[w(t)(\hat{\epsilon}_{\phi_{\text{LCN}}}(x_{t};y,t,I_{\text{cond}}(L,C))-\epsilon)\frac{\partial x}{\partial\Gamma(\beta(\cdot))}\right],$
@@
where $S$ denotes the object surface and $\epsilon$ is a small random 3D perturbation. We use $\lambda_{\text{recon}}=1000$ and $\lambda_{\text{reg}}=10$ to reweight the loss $\mathcal{L}_{\text{recon}}$ and $\mathcal{L}_{\text{reg}}$.
Scheduling the optimization. We warm up the optimization by rendering the four canonical views and applying $\mathcal{L}_{\text{recon}}$ for 50 iterations. We then add in iterations using the $\mathcal{L}_{\text{SDS}}$ loss and optimize over randomly chosen camera views and randomly selected lighting from a pre-defined set of environmental lighting maps. Specifically we alternate iterations between using $\mathcal{L}_{\text{SDS}}$ and $\mathcal{L}_{\text{recon}}$. In addition, for a quarter of the SDS iterations, we use the canonical views rather than randomly selecting the views. This ensures that the resulting texture does not overfit to the reference images corresponding to the canonical views. The warm-up iterations capture the large-scale structure of our texture and allow us to use relatively small noise levels ($t\leq 0.1$) in the SDS optimization. We sample the noise following a linearly decreasing schedule [17] with $t_{\text{max}}=0.1$ and $t_{\text{min}}=0.02$. We also adjust the conditioning strength $s$ of our LightControlNet in the SDS loss linearly from $1$ to 00 over these iterations so that LightControlNet is only lightly applied by the end of the optimization. We also experimented with a recent variant Variational Score Distillation [57], but did not observe notable improvement. We have experimentally found that we obtain high-quality textures after 400 total iterations of this optimization and this is far fewer iterations than other SDS-based texture generation techniques such as Fantasia3D [9] which requires 5000 iterations. More details are in the appendix.
Faster pipeline without relightability. Our two-stage pipeline is also compatible with off-the-shelf depth ControlNet and Stable Diffusion [44] as the backbone replacement of LightControlNet. Specifically, we can replace the LightControlNet in Stage 1 with a depth ControlNet that uses a depth rendering of the mesh as the conditioning image, and uses Stable Diffusion based SDS in Stage 2. In scenarios where texture relightability is not required, this variant offers an additional 2$\times$ speed-up (as shown in Table 1), since it eliminates the additional computation required by LightControlNet forward pass in the SDS optimization.